solve using N as the variable:
3 consecutive integers in which 4 times the sum of the first 2 is the same as 6 times the third?
please and thank you.

Answer by mbmodak(32)

let (n-1) , n ,(n+1) be three con integers
Now 4(n-1)+4n = 6(n+1) as per given condition
4n-4+4n = 6n+6
4n-4+4n-6n+4 = 6n+6-6n+4
8n-6n = 10
2n=10
n=10/2=5
the con integers are then (5-1) ,5, (5+1) or 4 , 5 ,6
[ 4*4+5*4=6*6 ]